# Scale Where It Matters: Training-Free Localized Scaling for Diffusion Models

### v2

@@ E.2 Scaling Comparison with Best-of-$N$
We evaluate how LoTTS scales with the total number of generated images $N$ and compare it against a pure global Best-of-$N$ baseline. In both cases, $N$ represents the total number of images generated per prompt. For Best-of-$N$, this is simply the number of globally sampled candidates. For LoTTS, however, the same budget $N$ is redistributed between the number of initial samples and the number of local refinements performed around each initial sample. For example, $N=1$ corresponds to pure global sampling, $N=3$ corresponds to $1$ initial sample with $k+1=3$ local candidates, $N=6$ corresponds to $2$ initial samples with $k+1=3$ refinements each, and $N=9$ corresponds to the default LoTTS setting of $3$ initial samples with $k+1=3$ refinements each.
Scaling behavior of LoTTS.   Tab. 8 shows that LoTTS exhibits smooth and consistent performance gains as the total number of generated images $N$ increases from 1 to 9. Even with extremely small budgets, such as $N=3$, LoTTS already achieves notable improvements over pure global sampling ($N=1$). For example, on SD2.1, HPS increases from 20.44 at $N=1$ to 21.12 at $N=3$, accompanied by an IR gain from 0.236 to 0.413. As $N$ continues increasing, these gains become more substantial: at $N=6$, SD2.1 reaches HPS 22.29 and IR 0.531; and at $N=9$, LoTTS achieves its strongest performance with HPS 24.52 and IR 0.680. A similar pattern appears in SDXL and FLUX. On SDXL, HPS improves from 23.44 ($N=1$) → 25.23 ($N=3$) → 26.21 ($N=6$) → 28.23 ($N=9$), with IR rising from 0.680 to 1.102. FLUX shows even stronger scaling: HPS increases from 29.34 ($N=1$) to 31.23 ($N=3$), then to 32.24 ($N=6$), and ultimately reaches 33.33 at $N=9$. The IR metric follows the same trend, improving from 1.038 to 1.605. These consistent improvements across all models and metrics indicate that each additional global starting point introduced by a larger budget $N$ provides new semantic modes for LoTTS to refine. Since every starting point is further expanded into multiple locally refined variants, LoTTS is able to generate a rich set of high-quality candidates around these diverse global modes.
@@
Scaling behavior of Best-of-$N$.   Tab. 9 shows that the Best-of-$N$ baseline consistently improves as the number of globally sampled images increases. This trend is expected: drawing more samples enlarges the global search space, increasing the likelihood of encountering rare but high-quality generations. As $N$ grows, the tails of the model’s output distribution are explored more thoroughly, naturally raising the probability of selecting an image with stronger aesthetics or semantic alignment. The numerical results reflect this behavior clearly. For SD2.1, HPS improves from 20.44 at N=1N{=}1 to 21.56 at N=9N{=}9, then to 22.21 at N=17N{=}17, ultimately reaching 24.23 at N=25N{=}25. SDXL follows a similar progression, increasing from 23.44 to 25.44, then to 26.02, and finally to 28.21 when N=30N{=}30. FLUX exhibits the same scaling pattern, with HPS rising from 29.34 at N=1N{=}1 to 30.23 at N=9N{=}9, then to 31.14 at N=14N{=}14, and ultimately reaching 33.32 at N=36N{=}36. The other metrics—AES, Pick, and IR—show comparable monotonic gains, though with diminishing increments as $N$ becomes large. Overall, these results confirm that Best-of-$N$ behaves like a global search procedure: as more samples are drawn, the chance of finding a superior candidate increases, leading to smooth and predictable scaling trends across models and evaluation metrics.
@@ F.1 Failure Analysis
Imperfect masks and unstable attention signals.   A major source of failure arises from attention-based masks missing true defects or selecting unstable high-frequency regions. Subtle artifacts, micro-textures, or weak structural inconsistencies may receive insufficient attention, leading to low recall, while noisy activations near boundaries often generate false positives. These behaviors correspond directly to the theoretical sensitivity to precision and recall discussed in Appendix G. In particular, insufficient recall violates the requirement in  equation G.3.1, preventing LoTTS from reaching enough defective patches to yield meaningful improvement. Conversely, poor precision inflates the false-positive harm term $\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$, which must remain smaller than $q\overline{\delta}$ for each selected patch to provide positive expected benefit, as formalized in  equation 22.
@@ F.2 Future Directions
Baseline beyond Best-of-$N$.   Although our current implementation adopts a Best-of-$N$ sampling scheme within the masked region, this strategy represents only the simplest form of local search. A more powerful alternative is to integrate global search algorithms—such as beam search—that branch along the denoising trajectory, but do so only within the masked area. Such mask-restricted branching preserves the locality and compute efficiency of LoTTS while enabling a substantially richer exploration of refinement hypotheses than Best-of-$N$. This opens the door to “global search applied locally,” where complex trajectory-level decisions are made without ever incurring the cost or harm of full-image resampling.
Early-stop mechanisms for low-potential cases.   Another practical extension is to incorporate an early-stop criterion that detects when an image has little potential for localized improvement. In many cases, the mask may identify only marginal or low-confidence regions, or the estimated patch-level quality signals may indicate that no meaningful refinement can be achieved within the masked area. Continuing localized resampling in such situations not only yields negligible benefit but also risks introducing unnecessary perturbations. An early-stop mechanism—triggered by low mask coverage, low defect confidence, or minimal improvement predicted from intermediate diffusion states—would allow the system to gracefully abstain from refinement when appropriate. This ensures that compute is spent only when LoTTS is theoretically capable of providing positive expected gain, further improving test-time efficiency and robustness.
Improving mask-generation strategies.   The current mask is constructed from relative internal inconsistencies, which can miss globally subtle or low-contrast defects. A more principled alternative is to incorporate absolute quality cues, such as no-reference IQA metrics, to better identify patches with genuine improvement potential. In addition, richer mask-generation strategies based on multi-signal fusion (e.g., trajectory disagreement, latent-gradient saliency) may significantly improve both precision and recall. By stabilizing the mask and grounding it in multiple complementary signals, these approaches can move the mask into the regime where LoTTS is theoretically guaranteed to yield positive expected gain ( equation 22), thereby enhancing robustness and overall refinement effectiveness.
Appendix G Theoretical Analysis
G.1 Problem Setup and Preliminaries
Overview.   Our theoretical analysis proceeds in three steps. (1) We first express global image quality through a patch-level lower bound, which allows local refinements to be analyzed independently. (2) We then characterize the stochastic behavior of the mask via precision and recall, which determines how often LoTTS touches defective or clean patches. (3) Combining these two components, we compute the expected gain of global and local TTS and derive the condition under which localized refinement is more compute-efficient.
@@ G.1 Problem Setup and Preliminaries
$S$ set of defective patches, with $|S|=s$
$\widehat{S}$ set of patches selected by LoTTS for resampling
@@
$w_{j}$ importance weight of patch $j$ in quality lower bound
@@
Table 10: Notation used in the theoretical analysis.
@@ G.1.1 Image Quality Decomposition
Before comparing global and local strategies, we first relate global quality $r(x)$ to patch-level scores. To lower bound global improvement by local improvements, we adopt a standard additive inequality:
@@
where $w_{j}\geq 0$ are weights reflecting importance of each patch. This allows us to measure global gain through weighted patch-level gains.
@@ G.1.2 Mask Selection Statistics
By the (operational) definition of precision $\pi=\frac{\mathbb{E}[\mathrm{TP}]}{\mathbb{E}[|\widehat{S}|]}$, we get:
$\mathbb{E}[|\widehat{S}|]=\frac{\mathbb{E}[\mathrm{TP}]}{\pi}=\frac{\rho\,s}{\pi}.$ (13)
Finally, using linearity of expectation and $\mathrm{FP}=|\widehat{S}|-\mathrm{TP}$,
$\mathbb{E}[\mathrm{FP}]=\mathbb{E}[|\widehat{S}|]-\mathbb{E}[\mathrm{TP}]=\frac{\rho\,s}{\pi}-\rho\,s=\rho\,s\Big(\tfrac{1}{\pi}-1\Big).$ (14)
@@ Discussion.
  • Why use $\pi=\frac{\mathbb{E}[\mathrm{TP}]}{\mathbb{E}[|\widehat{S}|]}$? Precision is often defined as $\mathbb{E}\!\left[\frac{\mathrm{TP}}{|\widehat{S}|}\mathbf{1}\{|\widehat{S}|>0\}\right]$. To avoid division-by-zero and keep algebra tractable, we adopt the ratio-of-expectations form, which is standard in compute-budget analyses. If one insists on the expectation-of-ratio definition, Jensen-type arguments yield the bound $\mathbb{E}[|\widehat{S}|]\geq\frac{\mathbb{E}[\mathrm{TP}]}{\pi}$, so our equalities become tight upper/lower bounds; the conclusions below only change by harmless inequalities.
@@
  • What randomness is averaged over? Expectations are taken over the randomness of mask construction (and, if applicable, sampling noise). In practice, the mask may be stochastic due to thresholding on noisy attention maps, variation across diffusion trajectories, or random perturbations used during test-time evaluation. No independence assumptions are required; linearity of expectation suffices.
@@ G.2 Main Theorem
G.2.1 Expected Gain per Resampling Step
We begin by analyzing the expected gain of a single resampling trial. Let $x$ denote the image before the update and $x^{+}$ the image after the update, and define
@@ G.2.1 Expected Gain per Resampling Step
To relate global quality to patch-level behavior, we use the standard additive lower bound:
@@
where $w_{j}\geq 0$ reflects the importance of patch $j$. This inequality allows us to lower-bound the global improvement through weighted patch-level changes:
$\Delta r\ \geq\ \sum_{j=1}^{M}w_{j}\big(r_{j}(x_{j}^{+})-r_{j}(x_{j})\big).$ (17)
For each defective patch $j\in S$, let $\delta_{j}\geq 0$ denote the expected improvement when the patch is successfully repaired; for each clean patch $j\notin S$, let $\gamma_{j}\geq 0$ denote the expected loss when the patch is harmed. Define the weighted averages
$\overline{\delta}:=\frac{1}{s}\sum_{j\in S}w_{j}\delta_{j},\hskip 28.80008pt\overline{\gamma}:=\frac{1}{M-s}\sum_{j\notin S}w_{j}\gamma_{j},$ (18)
where $s=|S|\ll M$ is the number of defective patches. These quantities do not require independence across patches; they capture only the expected per-patch effect of one resampling step.
Global resampling.   In global test-time scaling, all patches are resampled. Each defective patch $j\in S$ is repaired with probability $\theta_{g}$, and each clean patch $j\notin S$ is harmed with probability $h_{g}$. Aggregating over all patches gives the per-trial expected improvement:
$\mathbb{E}[\Delta r]_{\mathrm{global}}\ \geq\ s\,\theta_{g}\,\overline{\delta}\;-\;(M-s)\,h_{g}\,\overline{\gamma}.$ (19)
@@
Localized resampling (LoTTS).   In localized test-time scaling, only patches selected by the mask $\widehat{S}$ are resampled: potentially repairing true positives $S\cap\widehat{S}$ and harming false positives $\widehat{S}\setminus S$.
Let recall $\rho$ and precision $\pi$ be defined as in Lemma 1. Let $q$ be the probability of repairing a defective selected patch, and $h_{\ell}$ the probability of harming a selected clean patch. Using Lemma 1, the expected numbers of selected defective and clean patches are $\rho s$ and $\rho s(\tfrac{1}{\pi}-1)$, respectively, leading to
$\mathbb{E}[\Delta r]_{\mathrm{local}}\ \geq\ \rho\,s\,q\,\overline{\delta}\;-\;\rho\,s\Big(\tfrac{1}{\pi}-1\Big)\,h_{\ell}\,\overline{\gamma}.$ (20)
@@
G.2.2 Expected Gain per Compute Unit
To compare the two strategies under a fixed compute budget $B$, we normalize their per-trial gains by their respective costs $C_{g}$ and $C_{\ell}$. Since a total of $B/C_{g}$ global trials or $B/C_{\ell}$ local trials can be performed, we obtain:
@@ G.2.2 Expected Gain per Compute Unit
For a total compute budget $B$, let $C_{g}$ and $C_{\ell}$ denote the cost of a single global and localized trial, and let $\overline{\delta}$ and $\overline{\gamma}$ denote the average per–patch repair gain and harm penalty. Then the expected quality improvements of Global TTS and LoTTS satisfy: $\boxed{\begin{aligned} \mathbb{E}[\Delta r]_{\mathrm{global},B}&\;\geq\;\frac{B}{C_{g}}\Big[\,s\,\theta_{g}\,\overline{\delta}\;-\;(M-s)\,h_{g}\,\overline{\gamma}\,\Big],\\[3.0pt] \mathbb{E}[\Delta r]_{\mathrm{local},B}&\;\geq\;\frac{B}{C_{\ell}}\Big[\,\rho s\,q\,\overline{\delta}\;-\;\rho s\!\big(\tfrac{1}{\pi}-1\big)h_{\ell}\overline{\gamma}\,\Big].\end{aligned}}$
Interpretation.   Both methods yield an expected quality change equal to (average repair gain) minus (average harm penalty), scaled by the number of trials affordable under budget $B$.
These compute-normalized expressions provide the basis for the dominance condition analyzed in the next subsection.

### v3

@@ E.2 Scaling Comparison with Best-of-$N$
We evaluate how LoTTS scales with the total number of generated images $N$ and compare it against a pure global Best-of-$N$ baseline. In both cases, $N$ represents the total number of images generated per prompt. For Best-of-$N$, this is simply the number of globally sampled candidates. For LoTTS, however, the same budget $N$ is redistributed between the number of initial samples and the number of local refinements performed around each initial sample. For example, $N=1$ corresponds to pure global sampling, $N=3$ corresponds to $1$ initial sample producing $1+k=3$ total images (the original plus k=2k{=}2 local refinements), $N=6$ corresponds to $2$ initial samples each producing $1+k=3$ images, and $N=9$ corresponds to the default LoTTS setting of $3$ initial samples each producing $1+k=3$ images.
Scaling behavior of LoTTS. Tab. 12 shows that LoTTS exhibits overall consistent performance gains as the total number of generated images $N$ increases from 1 to 9. Even with extremely small budgets, such as $N=3$, LoTTS already achieves notable improvements over pure global sampling ($N=1$). For example, on SD2.1, HPS increases from 20.44 at $N=1$ to 21.12 at $N=3$, accompanied by an IR gain from 0.236 to 0.413. As $N$ continues increasing, these gains become more substantial: at $N=6$, SD2.1 reaches HPS 22.29 and IR 0.531; and at $N=9$, LoTTS achieves its strongest performance with HPS 24.52 and IR 0.680. A similar pattern appears in SDXL and FLUX. On SDXL, HPS improves from 23.44 ($N=1$) → 25.23 ($N=3$) → 26.21 ($N=6$) → 28.23 ($N=9$), with IR rising from 0.680 to 1.102. FLUX shows even stronger scaling: HPS increases from 29.34 ($N=1$) to 31.23 ($N=3$), then to 32.24 ($N=6$), and ultimately reaches 33.33 at $N=9$. The IR metric follows the same trend, improving from 1.038 to 1.605. These consistent improvements across all models and metrics indicate that each additional global starting point introduced by a larger budget $N$ provides new semantic modes for LoTTS to refine. Since every starting point is further expanded into multiple locally refined variants, LoTTS is able to generate a rich set of high-quality candidates around these diverse global modes.
@@
Scaling behavior of Best-of-$N$. Tab. 13 shows that the Best-of-$N$ baseline consistently improves as the number of globally sampled images increases. This trend is expected: drawing more samples enlarges the global search space, increasing the likelihood of encountering rare but high-quality generations. As $N$ grows, the tails of the model’s output distribution are explored more thoroughly, naturally raising the probability of selecting an image with stronger aesthetics or semantic alignment. The numerical results reflect this behavior clearly. For SD2.1, HPS improves from 20.44 at N=1N{=}1 to 21.56 at N=9N{=}9, then to 22.21 at N=17N{=}17, ultimately reaching 24.23 at N=25N{=}25. SDXL follows a similar progression, increasing from 23.44 to 24.54, then to 26.02, and finally to 28.21 when N=30N{=}30. FLUX exhibits the same scaling pattern, with HPS rising from 29.34 at N=1N{=}1 to 30.23 at N=9N{=}9, then to 31.14 at N=14N{=}14, and ultimately reaching 33.32 at N=36N{=}36. The other metrics, AES, Pick, and IR, show comparable monotonic gains, though with diminishing increments as $N$ becomes large. Overall, these results confirm that Best-of-$N$ behaves like a global search procedure: as more samples are drawn, the chance of finding a superior candidate increases, leading to smooth and predictable scaling trends across models and evaluation metrics.
@@ F.1 Failure Analysis
Imperfect masks and unstable attention signals. A major source of failure arises from attention-based masks missing true defects or selecting unstable high-frequency regions. Subtle artifacts, micro-textures, or weak structural inconsistencies may receive insufficient attention, leading to low recall, while noisy activations near boundaries often generate false positives. These behaviors correspond directly to the theoretical sensitivity to precision and recall discussed in Appendix G. In particular, insufficient recall violates the requirement in Eq. equation G.3, preventing LoTTS from reaching enough defective patches to yield meaningful improvement. Conversely, poor precision inflates the false-positive harm term $\rho(\tfrac{1}{\pi}-1)h_{\ell}\overline{\gamma}$, which must remain smaller than $q\overline{\delta}$ for each selected patch to provide positive expected benefit, as formalized in Eq. equation 23.
@@ F.2 Future Directions
Baseline beyond Best-of-$N$. Although our current implementation adopts a Best-of-$N$ sampling scheme within the masked region, this strategy represents only the simplest form of local search. A more powerful alternative is to integrate global search algorithms, such as beam search, that branch along the denoising trajectory, but do so only within the masked area. Such mask-restricted branching preserves the locality and compute efficiency of LoTTS while enabling a substantially richer exploration of refinement hypotheses than Best-of-$N$. This opens the door to “global search applied locally,” where complex trajectory-level decisions are made without ever incurring the cost or harm of full-image resampling.
Table 14: Notation used in the theoretical analysis.
@@
$\mathcal{D}$ set of defective patches, with $|\mathcal{D}|=s$
$\widehat{\mathcal{D}}$ set of patches selected by LoTTS for resampling
@@
$w_{j}$ importance weight of patch $j$ in quality decomposition
@@
Early-stop mechanisms for low-potential cases. Another practical extension is to incorporate an early-stop criterion that detects when an image has little potential for localized improvement. In many cases, the mask may identify only marginal or low-confidence regions, or the estimated patch-level quality signals may indicate that no meaningful refinement can be achieved within the masked area. Continuing localized resampling in such situations not only yields negligible benefit but also risks introducing unnecessary perturbations. An early-stop mechanism, triggered by low mask coverage, low defect confidence, or minimal improvement predicted from intermediate diffusion states, would allow the system to gracefully abstain from refinement when appropriate. This ensures that compute is spent only when LoTTS is theoretically capable of providing positive expected gain, further improving test-time efficiency and robustness.
Improving mask-generation strategies. The current mask is constructed from relative internal inconsistencies, which can miss globally subtle or low-contrast defects. A more principled alternative is to incorporate absolute quality cues, such as no-reference IQA metrics, to better identify patches with genuine improvement potential. In addition, richer mask-generation strategies based on multi-signal fusion (e.g., trajectory disagreement, latent-gradient saliency) may significantly improve both precision and recall. By stabilizing the mask and grounding it in multiple complementary signals, these approaches can move the mask into the regime where LoTTS is theoretically guaranteed to yield positive expected gain, as shown in Eq. equation 23, thereby enhancing robustness and overall refinement effectiveness.
Appendix G Theoretical Analysis
G.1 Problem Setup and Preliminaries
Overview. Our theoretical analysis proceeds in three steps. (1) We first decompose global image quality into patch-level contributions via an additive decomposition, which allows local refinements to be analyzed independently. (2) We then characterize the stochastic behavior of the mask via precision and recall, which determines how often LoTTS touches defective or clean patches. (3) Combining these two components, we compute the expected gain of global and local TTS and derive the condition under which localized refinement is more compute-efficient.
@@ G.1.1 Image Quality Decomposition
Before comparing global and local strategies, we first relate global quality $r(x)$ to patch-level scores. We adopt a standard additive decomposition assumption:
@@
where $w_{j}\geq 0$ are weights reflecting the importance of each patch. This decomposition holds exactly when the quality metric is patch-separable, as is the case for PSNR, patch-level LPIPS, and regional FID computed over local crops. It allows us to express global gain as a sum of weighted patch-level gains.
@@ G.1.2 Mask Selection Statistics
By the (operational) definition of precision $\pi=\frac{\mathbb{E}[\mathrm{TP}]}{\mathbb{E}[|\widehat{\mathcal{D}}|]}$, we get:
$\mathbb{E}[|\widehat{\mathcal{D}}|]=\frac{\mathbb{E}[\mathrm{TP}]}{\pi}=\frac{\rho\,s}{\pi}.$ (13)
Finally, using linearity of expectation and $\mathrm{FP}=|\widehat{\mathcal{D}}|-\mathrm{TP}$,
$\mathbb{E}[\mathrm{FP}]=\mathbb{E}[|\widehat{\mathcal{D}}|]-\mathbb{E}[\mathrm{TP}]=\frac{\rho\,s}{\pi}-\rho\,s=\rho\,s\bigl(\tfrac{1}{\pi}-1\bigr).$ (14)
@@ Discussion.
  • Why use $\pi=\frac{\mathbb{E}[\mathrm{TP}]}{\mathbb{E}[|\widehat{\mathcal{D}}|]}$?
    Precision is often defined as $\mathbb{E}\!\left[\frac{\mathrm{TP}}{|\widehat{\mathcal{D}}|}\mathbf{1}\{|\widehat{\mathcal{D}}|>0\}\right]$. To avoid division-by-zero and keep algebra tractable, we adopt the ratio-of-expectations form, which is standard in compute-budget analyses. If one insists on the expectation-of-ratio definition, Jensen-type arguments yield the bound $\mathbb{E}[|\widehat{\mathcal{D}}|]\geq\frac{\mathbb{E}[\mathrm{TP}]}{\pi}$, so our equalities become tight upper/lower bounds; the conclusions below only change by harmless inequalities.
  • Edge cases.
@@
  • What randomness is averaged over?
    Expectations are taken over the randomness of mask construction (and, if applicable, sampling noise). In practice, the mask may be stochastic due to thresholding on noisy attention maps, variation across diffusion trajectories, or random perturbations used during test-time evaluation. No independence assumptions are required; linearity of expectation suffices.
@@ G.2 Main Theorem
Expected Gain per Resampling Step. We begin by analyzing the expected gain of a single resampling trial. Let $x$ denote the image before the update and $x^{+}$ the image after the update, and define:
@@
To relate global quality to patch-level behavior, we use the additive decomposition Eq. equation 11:
@@
where $w_{j}\geq 0$ reflects the importance of patch $j$. Substituting the decomposition for both $x$ and $x^{+}$ yields:
$\Delta r\ =\ \sum_{j=1}^{M}w_{j}\big(r_{j}(x_{j}^{+})-r_{j}(x_{j})\big).$ (17)
For each defective patch $j\in\mathcal{D}$, let $\delta_{j}\geq 0$ denote the expected improvement when the patch is successfully repaired; for each clean patch $j\notin\mathcal{D}$, let $\gamma_{j}\geq 0$ denote the expected loss when the patch is harmed. Define the weighted averages:
$\overline{\delta}:=\frac{1}{s}\sum_{j\in\mathcal{D}}w_{j}\delta_{j},\qquad\overline{\gamma}:=\frac{1}{M-s}\sum_{j\notin\mathcal{D}}w_{j}\gamma_{j},$ (18)
where $s=|\mathcal{D}|\ll M$ is the number of defective patches. These quantities do not require independence across patches; they capture only the expected per-patch effect of one resampling step. For the local harm term derived below, we assume that false-positive patches are representative of clean patches in terms of $\gamma_{j}$; if mask-selected clean patches have systematically different harm, $\overline{\gamma}$ should be replaced by the corresponding conditional average.
Global resampling. In global test-time scaling, all patches are resampled. Each defective patch $j\in\mathcal{D}$ is repaired with probability $\theta_{g}$, and each clean patch $j\notin\mathcal{D}$ is harmed with probability $h_{g}$. Aggregating over all patches gives the per-trial expected improvement:
$\mathbb{E}[\Delta r]_{\mathrm{global}}\ =\ s\,\theta_{g}\,\overline{\delta}\;-\;(M-s)\,h_{g}\,\overline{\gamma}.$ (19)
@@
Localized resampling (LoTTS). In localized test-time scaling, only patches selected by the mask $\widehat{\mathcal{D}}$ are resampled: potentially repairing true positives $\mathcal{D}\cap\widehat{\mathcal{D}}$ and harming false positives $\widehat{\mathcal{D}}\setminus\mathcal{D}$. Let recall $\rho$ and precision $\pi$ be defined as in Lemma 1. Let $q$ be the probability of repairing a defective selected patch, and $h_{\ell}$ the probability of harming a selected clean patch. Using Lemma 1, the expected numbers of selected defective and clean patches are $\rho s$ and $\rho s(\tfrac{1}{\pi}-1)$, respectively, leading to:
$\mathbb{E}[\Delta r]_{\mathrm{local}}\ =\ \rho\,s\,q\,\overline{\delta}\;-\;\rho\,s\bigl(\tfrac{1}{\pi}-1\bigr)\,h_{\ell}\,\overline{\gamma}.$ (20)
@@
Expected Gain per Compute Unit. To compare the two strategies under a fixed compute budget $B$, we normalize their per-trial gains by their respective costs $C_{g}$ and $C_{\ell}$. Since a total of $B/C_{g}$ global trials or $B/C_{\ell}$ local trials can be performed, we obtain:
@@
For a total compute budget $B$, let $C_{g}$ and $C_{\ell}$ denote the cost of a single global and localized trial, and let $\overline{\delta}$ and $\overline{\gamma}$ denote the average per–patch repair gain and harm penalty. Under the additive quality decomposition Eq. equation 11, with each trial independently resampling from the same reference image [^1] In practice, all candidates are generated in parallel from the same base image and scored by a verifier that selects the best one (Best-of-NN). The linear scaling $\tfrac{B}{C}\Delta^{(1)}$ therefore serves as a compute-efficiency proxy: the strategy with higher $\Delta^{(1)}/C$ produces higher-quality candidates per unit of compute, and thus also yields a better Best-of-NN outcome under any fixed budget., the expected quality improvements of Global TTS and LoTTS are: $\mathbb{E}[\Delta r]_{\mathrm{global},B}$ $\;=\;\frac{B}{C_{g}}\bigl[\,s\,\theta_{g}\,\overline{\delta}\;-\;(M-s)\,h_{g}\,\overline{\gamma}\,\bigr],$ $\mathbb{E}[\Delta r]_{\mathrm{local},B}$ $\;=\;\frac{B}{C_{\ell}}\bigl[\,\rho s\,q\,\overline{\delta}\;-\;\rho s\!\big(\tfrac{1}{\pi}-1\big)h_{\ell}\overline{\gamma}\,\bigr].$
Interpretation. Both methods yield an expected quality change equal to (average repair gain) minus (average harm penalty), scaled by the number of trials affordable under budget $B$. These compute-normalized expressions provide the basis for the dominance condition analyzed in the next subsection.